OpenStudy (anonymous):

Hello, I'm having a little trouble with this one problem I'm on. It gives me vector r = {2t i + 4t^2 j} ft and it asks me to determine the radial component and the transverse component of the particle's velocity at t = 2 s. I've already found the radial component of the velocity, v_r = 16.0 ft/s. But I can't figure out how to find the transverse component. I know v_theta = r(deriv. of theta), but I can't figure out where I'm supposed to get theta from. I've tried a couple of things, but neither of them worked. Is there anyone online that can help me?

OpenStudy (anonymous):

@IrishBoy123 I see you're viewing my question. Might you be able/willing to help me?

OpenStudy (irishboy123):

sure

OpenStudy (anonymous):

OKay, cool. So like I said, I can't figure out what the question is trying to get me to do.

OpenStudy (irishboy123):

first of all we are working along the vector \(\vec r = <2t, 4t^2> = <4,16>_{t=2}\) which gives us \( \vec {\dot r} = <2,8t> = <2,16>_{t=2}\) is that agreed?

OpenStudy (anonymous):

Yes, that's correct.

OpenStudy (irishboy123):

you say \(v_r = 16\) how did you get that? because i see the need to resolve the components along \(\vec r\)

OpenStudy (anonymous):

That would be the magnitude of v_r, which is roughly 16.0 ft/s. MasteringEngineering said I was right, so I believe it is so. Although Mastering's answer is rounded a little different, as the actual answer I got was closer to 16.12 ft/s.

OpenStudy (irishboy123):

ok let me crunch some numbers then

OpenStudy (irishboy123):

mmm what you have calculated is \(|\vec{\dot r}| = \sqrt{2^2 + 16^2} = 16.12\) ie you have calculated the speed of the particle

OpenStudy (irishboy123):

if you resolve \(\vec{\dot r}\) along \(\vec r\), you get \(v_r = 16.000735\) is that closer to their answer?

OpenStudy (anonymous):

Yes. I found \[v _{r}\]. I am looking for \[v _{\theta}\], which is described as v_theta = r(dtheta/dt) (I don't know how to write it with the dot overhead on this thing).

OpenStudy (anonymous):

My problem is I don't know how to get theta, or if it's even relevant here, since everything seems to be in cartesian coordinates.

OpenStudy (irishboy123):

OK is my answer closer to the website's? we can deal with the rest in a minute

OpenStudy (irishboy123):



OpenStudy (anonymous):

Yes. to three sig figs, it is \[v _{r} = 16.0\]. But like I said, we already knew that one.

OpenStudy (irishboy123):

what they wany when they say radial velocity is this